OpenStudy (anonymous):

write each function in vertex form: 1) y=x^2-4x+6 2) y=4x^2+7x 3) y=-2x^2+8x+3

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\text{vertex form of parabola:}~~y=(a-h)^2+k\\~\\~\\ y=x^2-4x+6\\~\\~\\ \color{blue}{\text{this is little similar to squaring method of quadratic function}}\\~\\~\\ =(x^2-4x+4)-4+6\\~\\~\\ =(x-2)^2+2\\~\\~\\ \text{vertex(h,k)=(2,2)} }\end{align}\)

OpenStudy (mathmath333):

*quadratic equation

OpenStudy (mathmath333):

second equation is of the form (ax+b)^2=(ax)2+2axb+b^2 try to make adjustment to make it perfect square

OpenStudy (mathmath333):

*(ax+b)^2=(ax)^2+2axb+b^2